# SPHINX: Structural Prediction using Hypergraph Inference Network

### v1

@@ D.1 Full results on Particle Simulation
Golden standard represents the lower bound in terms of MSE for our model. Instead of predicting the hypergraph structure, this model uses the ground-truth connectivity. The oracle hypergraph is then fed into the hypergraph neural network for predicting the next timestep. The purpose of this model is to understand what is the target performance we can aim for, while taking into account the possible limitation of the hypergraph processor (AllDeepSets in our case).
Random structure follows the same pipeline used to train the Golden standard baseline. However, instead of the ground-truth connectivity, this model uses a randomly sampled hypergraph. The goal is to understand how much of our performance comes from having a strong decoder, compared to the advantage of predicting a real, suitable hypergraph structure.
@@
LSTM. Inspired by the baselines in [19], we designed an LSTM-based [55] model which receives all the nodes concatenated, and predicts, at each timestep, the future coordinates for all the nodes. While capable of capturing relationships between nodes, this model doesn’t have an inductive bias towards modelling higher-order interactions.
@@
Results.
Table 3 and Table 4 show the results of the ablation study on both the One-Triangle and Two-Triangles datasets. Compared to the model using either perturb structure (Random structure) or the one completely ignoring the higher-order connectivity (No structure), our model shows clearly advantage on both datasets, showing that learning to infer the hypergraphs is a critical step in the higher-order processing pipeline. Moreover, all the experiments show a clear advantage for the $k$-subset sampling compared to the classical Gumbel-Softmax, with the sequential prediction further increasing the performance. The LSTM baseline performs well on the next step prediction task, but the performance deteriorates severely on the medium ($5$ steps) and long-range prediction ($25$ steps). It is interesting to notice that, even if the model discovers the hypergraph structure almost perfectly (see Figure 4 in the main paper), there is still a gap in performance between our learnable SPHINX and the Golden standard model.
Figure 7: Experiments on Two-Triangles dataset investigating the importance of choosing the correct number of predicted hyperedges. Having less hyperedges than the real number (two) negatively impact the performance. However, when increasing the number of hyperedges above the true value, the performance does not deteriorate, indicating that the model is robust at discovering the hypergraph structure, as long as enough slots are provided.
D.2 Varying. number of hyperedges
@@ Results.
To establish to what extent the value we choose as the number of hyperedges impacts the performance, we conduct an ablation study by varying the number of slots predicted by our method. The results in Figure 7 show that picking a number of slots larger than the real one doesn’t harm the performance. Visual analysis of the learned hypergraph reveals that, when offering more slots than needed, the model tends to produce redundancy, associating the extra slots to hyperedges that were already discovered.

### v2

@@ F.1 Description of the models used in Particle Simulation
Random structure Instead of predicting the hypergraph structure, this model uses a randomly sampled hypergraph connectivity. The goal is to understand how much of our performance comes from having a strong decoder, compared to the advantage of predicting a real, suitable hypergraph structure.
@@
TDHNN and GroupNet We adapt the models proposed in  (Zhou et al., 2023) and (Xu et al., 2022) respectively to predict the group of particles that are interacting in the Particle Simulation dataset. To ensure a fair comparison we are using the same decoder in all experiments and only modify the encoder.
F.2 Varying. number of hyperedges
@@
To establish to what extent the value we choose as the number of hyperedges impacts the performance, we conduct an ablation study by varying the number of slots predicted by our method. The results in Figure 5 show that picking a number of slots larger than the real one (2) doesn’t harm the performance. Visual analysis of the learned hypergraph reveals that, when offering more slots than needed, the model tends to produce redundancy, associating the extra slots to hyperedges that were already discovered. To better understand this behaviour, we measured the average number of distinct hyperedges predicted when providing more slots than required. The results are show in Table 5.
Table 5: Ablation study providing more slots than the real number. Unique Hyperedges denotes the number of unique predicted hyperedges while MSE denotes the final performance. The real number of higher-order interactions for this experiment is 2.
Slots Unique Hyperedges MSE
1 slot 1.00 0.00007
2 slots 1.96 0.00002
3 slots 2.15 0.000017
4 slots 2.14 0.000007
We believe that this is a very interesting result, as it shows that when equipped with more slots than necessary the model learns some form of redundancy (by predicting the same hyperedge multiple times) instead of hallucinating fake relations.